# Fourier Circuits in Neural Networks: Unlocking the Potential of Large Language Models in Mathematical Reasoning and Modular Arithmetic

### v1

@@
In the evolving landscape of machine learning, a pivotal challenge lies in deciphering the internal representations harnessed by neural networks and Transformers. Building on recent progress toward comprehending how networks execute distinct target functions, our study embarks on an exploration of the underlying reasons behind networks adopting specific computational strategies. We direct our focus to the complex algebraic learning task of modular addition involving $k$ inputs. Our research presents a thorough analytical characterization of the features learned by stylized one-hidden layer neural networks and one-layer Transformers in addressing this task.
A cornerstone of our theoretical framework is the elucidation of how the principle of margin maximization shapes the features adopted by one-hidden layer neural networks. Let $p$ denote the modulus, $D_{p}$ denote the dataset of modular arithmetic with $k$ inputs and $m$ denote the network width. We demonstrate that a neuron count of $m\geq 2^{2k-2}\cdot(p-1)$, these networks attain a maximum $L_{2,k+1}$-margin on the dataset $D_{p}$. Furthermore, we establish that each hidden-layer neuron aligns with a specific Fourier spectrum, integral to solving modular addition problems.
By correlating our findings with the empirical observations of similar studies, we contribute to a deeper comprehension of the intrinsic computational mechanisms of neural networks. Furthermore, we observe similar computational mechanisms in the attention matrix of the Transformer. This research stands as a significant stride in unraveling their operation complexities, particularly in the realm of complex algebraic tasks.
@@ 1 Introduction
The field of artificial intelligence has experienced a significant transformation with the development of large language models (LLMs), particularly through the introduction of the Transformer architecture. This advancement has revolutionized approaches to challenging tasks in natural language processing, notably in machine translation [67, 27] and text generation [52]. Consequently, models e.g., BERT [19], PaLM [16], LLaMA 2[86], ChatGPT [64], GPT4 [65] and so on, have become predominant in NLP.
@@
Central to this study is the question of how these advanced models transcend mere pattern recognition to engage in what appears to be logical reasoning and problem-solving. This inquiry is not purely academic; it probes the core of “understanding” in artificial intelligence. While LLMs, such as GPT, demonstrate remarkable proficiency in human-like text generation, their capability in comprehending and processing mathematical logic is a topic of considerable debate.
@@
This line of investigation is crucial, given AI’s potential to extend beyond text generation into deeper comprehension of complex subjects. Mathematics, often seen as the universal language, presents a uniquely challenging domain for these models [95]. Our research aims to determine if Transformers, noted for their NLP efficiency, can also demonstrate an intrinsic understanding of mathematical operations and reasoning.
@@
A recent surprising study of mathematical operations learning, [66] train Transformers on small algorithmic datasets, e.g., $a_{1}+a_{2}\mod p$ and we let $p$ be a prime number, and show the “grokking” phenomenon, where models abruptly transition from bad generalization to perfect generalization after a large number of training steps. Nascent studies, such as those by [60], empirically reveal that Transformers can solve modular addition using Fourier-based circuits. They found that the Transformers trained by Stochastic Gradient Descent (SGD) not only reliably compute $a_{1}+a_{2}\mod p$, but also that the networks consistently employ a specific geometric algorithm. This algorithm, which involves composing integer rotations around a circle, indicates an inherent comprehension of modular arithmetic within the network’s architecture. The algorithm relies on this identity: for any $a_{1},a_{2}$ and $\zeta\in\mathbb{Z}_{p}\backslash\{0\}$, the following two quantities are equivalent
  • $(a_{1}+a_{2})\bmod p$
  • $\underset{c\in\mathbb{Z}_{p}}{\arg\max}\{\cos({2\pi\zeta(a_{1}+a_{2}-c)}/{p})\}$
[60] further show that the attention and MLP module in the Transformer imbues the neurons with Fourier circuit-like properties.
@@
To study why networks arrive at Fourier-based circuits computational strategies, [55] theoretically study one-hidden layer neural network learning on two inputs modular addition task and certify that the trained networks will execute modular addition by employing Fourier features aligning closely with the previous empirical observations. However, the question remains whether neural networks can solve more complicated mathematical problems.
@@
Inspired by recent developments in mechanistic interpretability [62, 23, 22] and the study of inductive biases [72, 88] in neural networks, we extend our research to modular addition with three or more ($k$) inputs.
$(a_{1}+\dots+a_{k})\bmod p.$ (1)
@@
  • Expansion of Input for Cyclic Groups Problem: We extend the input parameter range for the cyclic groups problem from a binary set to $k$-element sets.
  • Network’s Maximum Margin: Let $L_{a,b}$ be defined in Definition 3.3. On the modular dataset $D_{p}$, we demonstrate that the maximum $L_{2,k+1}$-margin of a network is:
@@
  • Neuron Count in One-Hidden-Layer Networks: We propose that in a general case, a one-hidden-layer network having $m\geq 2^{2k-2}\cdot(p-1)$ neurons can achieve the maximum $L_{2,k+1}$-margin solution. This ensures the network’s capability to effectively solve the cyclic groups problem in a Fourier-based method, a finding corroborated by our experimental data.
  • Empirical Validation of Theoretical Findings: We validate our theoretical finding that when $m\geq 2^{2k-2}\cdot(p-1)$, for each spectrum $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$, there exists a hidden-neuron utilizes this spectrum.
  • Similar Empirical Findings in One-Layer Transformer: We also have a similar observation in one-layer Transformer learning modular addition involving $k$ inputs. For the $2$-dimensional matrix $W_{K}W_{Q}$, where $W_{K},W_{Q}$ denotes the key and query matrix, it shows the superposition of two cosine waveforms in each dimension, each characterized by distinct frequencies.
  • Grokking under Different $k$: We observe that as $k$ increases, the grokking phenomenon becomes weak.
@@ 2 Related Work
Max Margin Solutions in Neural Networks.
[3] demonstrated that neurons in a one-hidden-layer ReLU network align with clauses in max margin solutions for read-once DNFs, employing a unique proof technique involving the construction of perturbed networks. [55] utilize max-min duality to certify maximum-margin solutions. Further, extensive research in the domain of margin maximization in neural networks, including works by [29, 72, 28, 89, 49, 41, 59, 10, 42, 51, 25, 26] and more, has highlighted the implicit bias towards margin maximization inherent in neural network optimization. They provide a foundational understanding of the dynamics of neural networks and their inclination towards maximizing margins under various conditions and architectures.
Algebraic Tasks Learning Mechanism Interpretability.
The study of neural networks trained on algebraic tasks has been pivotal in shedding light on their training dynamics and inductive biases. Notable contributions include the work of [66, 12] on modular addition and subsequent follow-up studies, investigations into learning parities [21, 4, 80, 82, 81], and research into algorithmic reasoning capabilities [71, 32, 45, 20]. The field of mechanistic interpretability, focusing on the analysis of internal representations in neural networks, has also seen significant advancements through the works of [13, 63] and others.
Grokking and Emergent Ability.
The phenomenon known as “grokking” was initially identified by [66] and is believed to be a way of studying the emerging abilities of LLM [91]. This research observed a unique trend in two-layer transformer models engaged in algorithmic tasks, where there was a significant increase in test accuracy, surprisingly occurring well after these models had reached perfect accuracy in their training phase. In [56], it was hypothesized that this might be the result of the SGD process that resembles a random path along what is termed the optimal manifold. Adding to this, [60] aligns with the findings of [5], indicating a steady advancement of networks towards algorithms that are better at generalization. [48, 94, 47] developed smaller-scale examples of grokking and utilized these to map out phase diagrams, delineating multiple distinct learning stages. Furthermore, [85, 58] suggested the possibility of grokking occurring naturally, even in the absence of explicit regularization. They attributed this to an optimization quirk they termed the slingshot mechanism, which might inadvertently act as a regularizing factor.
@@
To calculate Fourier transform there are two main methodologies: one uses carefully chosen samples through hashing functions (referenced in works like [30, 33, 34, 37, 36, 43, 44]) to achieve sublinear sample complexity and running time, while the other uses random samples (as discussed in [18, 69, 8, 35, 61]) with sublinear sample complexity but nearly linear running time. There are many other works studying Fourier transform [68, 57, 75, 39, 31, 53, 15, 77, 14, 78, 17, 83].
@@ 3 Problem Setup
Part of our notations are following [55]. The setting of our one-hidden-layer neural network is in Section 3.1. In Section 3.2, we define the margin of the neuron network. We introduce a lemma that connects the training neuron network to solving the maximum-margin problem in Section 3.3.
@@ 3.1 Data and Network Setup
We denote $\mathbb{Z}_{p}$ as the modular group on $p$ integers, e.g., $\mathbb{Z}_{p}=[p]$, where $p>2$ is a given prime number. We denote $\mathcal{X}:=\mathbb{Z}_{p}^{k}$ as the input space and $\mathcal{Y}:=\mathbb{Z}_{p}$ as the output space. We denote $D_{p}:=\{((a_{1},\dots,a_{k}),\sum_{i\in[k]}a_{i}):a_{1},\dots,a_{k}\in\mathbb{Z}_{p}\}$ as the modular dataset.
Suppose we have a $L_{2,k+1}$ (matrix) norm $\|\cdot\|_{2,k+1}$ (Definition 3.3) and a class of parameterized functions $\{f(\theta,\cdot)\mid\theta\in\mathbb{R}^{U}\}$, where $f:\mathbb{R}^{U}\times\mathcal{X}\rightarrow\mathbb{R}^{\mathcal{Y}}$ and a matrix set $\Theta:=\{\|\theta\|_{2,k+1}\leq 1\}$. We consider single-hidden layer neural networks with polynomial activation functions and without biases. With parameters $\theta\in\Theta$, we denote $f(\theta,x)$ as the network’s output for an input $x$. For one-hidden layer networks, we denote $f$ as:
$f(\theta,x):=\sum_{i=1}^{m}\phi(\theta_{i},x),$
where $\theta:=\{\theta_{1},\ldots,\theta_{m}\}$. Here, we denote $\phi$ as one neuron and $\theta_{i}\in\Omega$ are the corresponding weights in that neuron, where $\Omega$ is a vector set. Let $\Omega^{\prime}$ be a subset of $\Omega$. For every $i\in[m]$, when either there exists $\alpha_{i}>0$ such that $\alpha_{i}\theta_{i}\in\Omega^{\prime}$ or $\theta_{i}=0$, then we say the parameter set $\theta=\{\theta_{1},\ldots,\theta_{m}\}$ has directional support on $\Omega^{\prime}$. A single neuron is represented as:
$\phi(\{u_{1},\dots,u_{k},w\},x_{1},\dots,x_{k}):=~{}(u_{1}^{\top}x_{1}+\dots+u_{k}^{\top}x_{k})^{k}w,$
where $u_{1},\dots,u_{k},w\in\mathbb{R}^{p}$ are the neuron’s weights, and $x_{1},\dots,x_{k}\in\mathbb{R}^{p}$ are the network inputs. Inputs $x_{1},\dots,x_{k}$ are one-hot vectors representing group elements. For input elements $(a_{1},\dots,a_{k})$, a neuron simplifies to
$\phi(\{u_{1},\dots,u_{k},w\},a_{1},\dots,a_{k})=(u_{1}(a_{1})+\dots+u_{k}(a_{k}))^{k}w,$
with $u_{1}(a_{1})$ being the $a_{1}$-th component of $u_{1}$. The output of the neuron is in $p$-dimension for cross-entropy loss. With $\theta=\{u_{i,1},\dots,u_{i,k},w_{i}\}_{i=1}^{m},$ the network is denoted as:
$~{}f(\theta,a_{1},\dots,a_{k}):=~{}\sum_{i=1}^{m}\phi(\{u_{i,1},\dots,u_{i,k},w_{i}\},a_{1},\dots.a_{k}).$ (2)
We define our regularized training objective function.
@@
3.2 Definition and Notation
We define the vector norm and matrix norm as the following.
Definition 3.2 ($L_{b}$ (vector) norm).
Definition 3.3 ($L_{a,b}$ (matrix) norm).
@@ 3.2 Definition and Notation
We denote $g:\mathbb{R}^{U}\times\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ as the margin function, where for given $(x,y)\in D_{p}$,
$g(\theta,x,y):=f(\theta,x)[y]-\max_{y^{\prime}\in\mathcal{Y}\backslash\{y\}}f(\theta,x)[y^{\prime}].$
@@
The margin for a given dataset $D_{p}$ is denoted as $h:\mathbb{R}^{U}\rightarrow\mathbb{R}$ where
$h(\theta):=\min_{(x,y)\in D_{p}}g(\theta,x,y).$
@@
The minimum of the regularized objective is denoted as $\theta_{\lambda}\in\arg\min_{\theta\in\mathbb{R}^{U}}\mathcal{L}_{\lambda}(\theta)$. We define the normalized margin of $\theta_{\lambda}$ as $\gamma_{\lambda}:=h(\theta_{\lambda}/\|\theta_{\lambda}\|_{2,k+1})$. We define the maximum normalized margin as $\gamma^{*}:=\max_{\theta\in\Theta}h(\theta)$.
Let $\mathcal{P}(D_{p})$ denote as a set containing any distributions over the dataset $D_{p}$. We see that $\gamma^{*}$ can be rewritten as
$\gamma^{*}=$ $~{}\max_{\theta\in\Theta}h(\theta)$
$=$ $~{}\max_{\theta\in\Theta}\min_{(x,y)\in D_{p}}g(\theta,x,y)$
$=$ $~{}\max_{\theta\in\Theta}\min_{q\in\mathcal{P}(D_{p})}\underset{(x,y)\sim q}{\mathbb{E}}[g(\theta,x,y)],$ (3)
@@
We define a pair $(\theta^{*},q^{*})$ when satisfying
$q^{*}$ $\in\underset{q\in\mathcal{P}(D_{p})}{\arg\min}\underset{(x,y)\sim q}{\mathbb{E}}[g(\theta^{*},x,y)]$ (4)
$\theta^{*}$ $\in\underset{\theta\in\Theta}{\arg\min}\underset{(x,y)\sim q^{*}}{\mathbb{E}}[g(\theta,x,y)].$ (5)
This means that $q^{*}$ is among the entities that minimize the expected margin based on $\theta^{*}$, while $\theta^{*}$ is among the entities that maximize the expected margin relative to $q^{*}$. The max-min inequality, as referenced in [9], indicates that presenting such a duo adequately proves $\theta^{*}$ to be a maximum margin solution.
Recall that there is a “max” operation in Definition 3.4, which makes the swapping of expectation and summation infeasible, which means that the expected network margin cannot be broken down into the expected margins of individual neurons. To tackle this problem, the class-weighted margin is proposed. Let $\tau:D_{p}\rightarrow\Delta(\mathcal{Y})$ allocate weights to incorrect labels for every data point. Given $(x,y)$ in $D_{p}$ and for any $y^{\prime}\in\mathcal{Y}$, we have $\tau(x,y)[y^{\prime}]\geq 0$ and $\sum_{y^{\prime}\in\mathcal{Y}\backslash\{y\}}\tau(x,y)[y^{\prime}]=1$. Then, we denote $g^{\prime}$ as the following to solve the issue.
@@
$~{}\underset{(x,y)}{\operatorname*{{\mathbb{E}}}}[g^{\prime}(\theta,x,y)]=~{}\sum_{i=1}^{m}\underset{(x,y)}{\operatorname*{{\mathbb{E}}}}[\phi(\theta_{i},x)[y]-\sum_{y^{\prime}\in\mathcal{Y}\backslash\{y\}}\tau(x,y)[y^{\prime}]\phi(\theta_{i},x)[y^{\prime}]].$
3.3 Preliminary

### v2

@@
In the evolving landscape of machine learning, a pivotal challenge lies in deciphering the internal representations harnessed by neural networks and Transformers. Building on recent progress toward comprehending how networks execute distinct target functions, our study embarks on an exploration of the underlying reasons behind networks adopting specific computational strategies. We direct our focus to the complex algebraic learning task of modular addition involving $k$ inputs. Our research presents a thorough analytical characterization of the features learned by stylized one-hidden layer neural networks and one-layer Transformers in addressing this task. A cornerstone of our theoretical framework is the elucidation of how the principle of margin maximization shapes the features adopted by one-hidden layer neural networks. Let $p$ denote the modulus, $D_{p}$ denote the dataset of modular arithmetic with $k$ inputs and $m$ denote the network width. We demonstrate that a neuron count of $m\geq 2^{2k-2}\cdot(p-1)$, these networks attain a maximum $L_{2,k+1}$-margin on the dataset $D_{p}$. Furthermore, we establish that each hidden-layer neuron aligns with a specific Fourier spectrum, integral to solving modular addition problems. By correlating our findings with the empirical observations of similar studies, we contribute to a deeper comprehension of the intrinsic computational mechanisms of neural networks. Furthermore, we observe similar computational mechanisms in the attention matrix of the one-layer Transformer. This research stands as a significant stride in unraveling their operation complexities, particularly in the realm of complex algebraic tasks.
@@ 1 Introduction
The field of artificial intelligence has experienced a significant transformation with the development of large language models (LLMs), particularly through the introduction of the Transformer architecture [108]. This advancement has revolutionized approaches to challenging tasks in natural language processing, notably in machine translation [78, 29] and text generation [60]. Consequently, models e.g., BERT [21], PaLM [18], Mistral [47], Llama [101], Llama2 [104], Llama3 [1], Gemini [98], Gemma [103], Claude3 [2], ChatGPT [75], GPT4 [76] and so on, have become predominant in NLP.
@@
Central to this study is the question of how these advanced models transcend mere pattern recognition to engage in what appears to be logical reasoning and problem-solving. This inquiry is not purely academic; it probes the core of “understanding” in artificial intelligence. While LLMs, such as Claude3 and GPT4, demonstrate remarkable proficiency in human-like text generation, their capability to comprehend and process mathematical logic is a topic of considerable debate.
@@
This line of investigation is crucial, given AI’s potential to extend beyond text generation into deeper comprehension of complex subjects. Mathematics, often seen as the universal language, presents a uniquely challenging domain for these models [116]. Our research aims to determine whether Transformers with attention, noted for their NLP efficiency, can also demonstrate an intrinsic understanding of mathematical operations and reasoning.
@@
In a recent surprising study of mathematical operations learning, [77] train Transformers on small algorithmic datasets, e.g., $a_{1}+a_{2}\mod p$ and we let $p$ be a prime number, and show the “grokking” phenomenon, where models abruptly transition from bad generalization to perfect generalization after a large number of training steps. Nascent studies, such as those by [70], empirically reveal that Transformers can solve modular addition using Fourier-based circuits. They found that the Transformers trained by Stochastic Gradient Descent (SGD) not only reliably compute $a_{1}+a_{2}\mod p$, but also that the networks consistently employ a specific geometric algorithm. This algorithm, which involves composing integer rotations around a circle, indicates an inherent comprehension of modular arithmetic within the network’s architecture. The algorithm relies on this identity: for any $a_{1},a_{2}$ and $\zeta\in\mathbb{Z}_{p}\backslash\{0\}$, the following two quantities are equivalent
$(a_{1}+a_{2})\bmod p=\underset{c\in\mathbb{Z}_{p}}{\arg\max}\{\cos({2\pi\zeta(a_{1}+a_{2}-c)}/{p})\}.$
@@
[70] further show that the attention and MLP module in the Transformer imbues the neurons with Fourier circuit-like properties. To study why networks arrive at Fourier-based circuits computational strategies, [64] theoretically study one-hidden layer neural network learning on two inputs modular addition task and certify that the trained networks will execute modular addition by employing Fourier features aligning closely with the previous empirical observations. However, the question remains whether neural networks can solve more complicated mathematical problems.
@@
Inspired by recent developments in mechanistic interpretability [73, 25, 24] and the study of inductive biases [85, 106] in neural networks, we extend our research to modular addition with more ($k$) inputs.
$(a_{1}+a_{2}+\dots+a_{k})\bmod p.$ (1)
@@
  • Expansion of Input for Modular Addition Problem: We extend the input parameter range for the modular addition problem from a binary set to $k$-element sets.
  • Network’s Maximum Margin: For $p$-modular addition of $k$ inputs, we give the closed form of the maximum margin of a network (Lemma 4.2):
@@
  • Neuron Count in One-Hidden-Layer Networks: We propose that in a general case, a one-hidden-layer network having $m\geq 2^{2k-2}\cdot(p-1)$ neurons can achieve the maximum $L_{2,k+1}$-margin solution. This ensures the network’s capability to effectively solve the modular addition problem in a Fourier-based method (Theorem 4.1).
  • Empirical Validation of Theoretical Findings: We validate our theoretical finding that when $m\geq 2^{2k-2}\cdot(p-1)$, for each spectrum $\zeta\in\{1,\ldots,\frac{p-1}{2}\}$, there exists a hidden-neuron utilizes this spectrum, strong empirical support of our analysis. (Figure 1 and Figure 2).
  • Similar Findings in Transformer: We have a similar observation in one-layer Transformer learning modular addition involving $k$ inputs. For the $2$-dimensional matrix $W_{K}W_{Q}$, where $W_{K},W_{Q}$ denotes the key and query matrix, it shows the superposition of two cosine waveforms in each dimension, each characterized by distinct frequencies (Figure 3).
  • Grokking under Different $k$: We observe that as $k$ increases, the grokking phenomenon becomes weaker, as predicted by our analysis (Figure 4).
@@ 2 Related Work
Max Margin Solutions in Neural Networks. [5] demonstrated that neurons in a one-hidden-layer ReLU network align with clauses in max margin solutions for read-once DNFs, employing a unique proof technique involving the construction of perturbed networks. [64] utilize max-min duality to certify maximum-margin solutions. Further, extensive research in the domain of margin maximization in neural networks, including works by [33, 85, 32, 109, 57, 48, 69, 12, 49, 59, 27, 28, 87, 30] and more, has highlighted the implicit bias towards margin maximization inherent in neural network optimization. They provide a foundational understanding of the dynamics of neural networks and their inclination towards maximizing margins under various conditions and architectures.
Algebraic Tasks Learning Mechanism Interpretability. The study of neural networks trained on algebraic tasks has been pivotal in shedding light on their training dynamics and inductive biases. Notable contributions include the work of [77, 35, 80] on modular addition and subsequent follow-up studies, investigations into learning parities [23, 6, 93, 95, 94, 83, 118, 112], and research into algorithmic reasoning capabilities [84, 37, 53, 63, 22, 14, 96, 71, 117, 100, 40]. The field of mechanistic interpretability, focusing on the analysis of internal representations in neural networks, has also seen significant advancements through the works of [15, 74, 68, 81, 107] and others.
Grokking and Emergent Ability. The phenomenon known as “grokking” was initially identified by [77] and is believed to be a way of studying the emerging abilities of LLM [111]. This research observed a unique trend in two-layer transformer models engaged in algorithmic tasks, where there was a significant increase in test accuracy, surprisingly occurring well after these models had reached perfect accuracy in their training phase. In [65], it was hypothesized that this might be the result of the SGD process that resembles a random path along what is termed the optimal manifold. Adding to this, [70] aligns with the findings of [7], indicating a steady advancement of networks towards algorithms that are better at generalization. [56, 115, 55] developed smaller-scale examples of grokking and utilized these to map out phase diagrams, delineating multiple distinct learning stages. Furthermore, [102, 67] suggested the possibility of grokking occurring naturally, even in the absence of explicit regularization. They attributed this to an optimization quirk they termed the slingshot mechanism, which might inadvertently act as a regularizing factor.
@@
To calculate Fourier transform there are two main methodologies: one uses carefully chosen samples through hashing functions (referenced in works like [34, 38, 39, 43, 42, 50, 51]) to achieve sublinear sample complexity and running time, while the other uses random samples (as discussed in [20, 82, 10, 41, 72]) with sublinear sample complexity but nearly linear running time. There are many other works studying Fourier transform [79, 66, 88, 45, 36, 61, 17, 90, 16, 91, 19, 97, 31].
@@ 3.1 Data and Network Setup
Data. Following [64], let $\mathbb{Z}_{p}=[p]$ denote the modular group on $p$ integers, where $p>2$ is a given prime number. The input space is $\mathcal{X}:=\mathbb{Z}_{p}^{k}$ for some integer $k$, and the output space is $\mathcal{Y}:=\mathbb{Z}_{p}$. Then an input data point is $a=(a_{1},\dots,a_{k})$ with $a_{i}\in\mathbb{Z}_{p}$. When clear from context, we also let $x_{i}\in\{0,1\}^{p}$ be the one-hot encoding of $a_{i}$, and let $x=(x_{1},\dots,x_{k})$ denote the input point.
Network. We consider single-hidden layer neural networks with polynomial activation functions:
$f(\theta,x):=\sum_{i=1}^{m}\phi(\theta_{i},x),\quad\quad\phi(\theta_{i},x):=(u_{i,1}^{\top}x_{1}+\dots+u_{i,k}^{\top}x_{k})^{k}w_{i},$ (2)
where $\theta:=\{\theta_{1},\ldots,\theta_{m}\}$,  $\phi(\theta_{i},x)$ is one neuron, and $\theta_{i}:=\{u_{i,1},\dots,u_{i,k},w_{i}\}$ are the parameters of the neuron, with $u_{i,1},\dots,u_{i,k},w_{i}\in\mathbb{R}^{p}$. We use polynomial activation functions as the homogeneous requirement in Lemma 3.7 and easy sum-to-product identities calculation in Fourier analysis. Using the notation $a$ instead of the one-hot encodings $x$, we can also write:
$f(\theta,a):=\sum_{i=1}^{m}\phi(\theta_{i},a),\quad\quad\phi(\theta_{i},a):=(u_{i,1}(a_{1})+\dots+u_{i,k}(a_{k}))^{k}w_{i},$
where with $u_{i,j}(a_{j})$ being the $a_{j}$-th component of $u_{i,j}$. We consider the parameter set:
$\Theta:=\{\|\theta\|_{2,k+1}\leq 1\},\text{~{}~{}where~{}}\|\theta\|_{2,k+1}:=(\sum_{i=1}^{m}\|\theta_{i}\|_{2}^{k+1})^{1\over k+1},\text{~{}and~{}}\|\theta_{i}\|_{2}:=(\sum_{j=1}^{k}\|u_{i,j}\|^{2}_{2}+\|w_{i}\|_{2}^{2})^{1\over 2}.$
Here $\|\theta\|_{2,k+1}$ is the $L_{2,k+1}$ matrix norm of $\theta$ (Definition C.2), and $\|\theta_{i}\|_{2}$ is the $L_{2}$ vector norm of the concatenated vector of the parameters in $\theta_{i}$. The training objective over $\Theta$ is then as follows.
@@
Given a dataset $D_{p}$ and the cross-entropy loss $l$, the regularized training objective is:
@@
3.2 Margins of the Neural Networks
Now, we define the margin for a data point and the margin for a whole dataset.
@@ 3.2 Margins of the Neural Networks
We denote $g:\mathbb{R}^{U}\times\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ as the margin function, where for given $(x,y)\in D_{p}$, $g(\theta,x,y):=f(\theta,x)[y]-\max_{y^{\prime}\in\mathcal{Y}\backslash\{y\}}f(\theta,x)[y^{\prime}].$
@@
The margin for a given dataset $D_{p}$ is denoted as $h:\mathbb{R}^{U}\rightarrow\mathbb{R}$ where $h(\theta):=\min_{(x,y)\in D_{p}}g(\theta,x,y).$
@@
The minimum of the regularized objective is denoted as $\theta_{\lambda}\in\arg\min_{\theta\in\mathbb{R}^{U}}\mathcal{L}_{\lambda}(\theta)$. We define the normalized margin of $\theta_{\lambda}$ as $\gamma_{\lambda}:=h(\theta_{\lambda}/\|\theta_{\lambda}\|_{2,k+1})$. We define the maximum normalized margin as $\gamma^{*}:=\max_{\theta\in\Theta}h(\theta)$, where $\Theta=\{\|\theta\|_{2,k+1}\leq 1\}$.
Let $\mathcal{P}(D_{p})$ denote as a set containing any distributions over $D_{p}$. Then $\gamma^{*}$ can be rewritten as
$\gamma^{*}=$ $~{}\max_{\theta\in\Theta}h(\theta)=~{}\max_{\theta\in\Theta}\min_{(x,y)\in D_{p}}g(\theta,x,y)=~{}\max_{\theta\in\Theta}\min_{q\in\mathcal{P}(D_{p})}\underset{(x,y)\sim q}{\mathbb{E}}[g(\theta,x,y)],$ (3)
@@
Now, we introduce an important concept of a duality stationary pair $(\theta^{*},q^{*})$.
@@
We define a stationary pair $(\theta^{*},q^{*})$ when satisfying
$q^{*}$ $\in\underset{q\in\mathcal{P}(D_{p})}{\arg\min}\underset{(x,y)\sim q}{\mathbb{E}}[g(\theta^{*},x,y)],\quad\quad\theta^{*}\in\underset{\theta\in\Theta}{\arg\min}\underset{(x,y)\sim q^{*}}{\mathbb{E}}[g(\theta,x,y)].$ (4)
This means that $q^{*}$ is a distribution that minimizes the expected margin based on $\theta^{*}$, and simultaneously, $\theta^{*}$ is a solution that maximizes the expected margin relative to $q^{*}$. The max-min inequality [11] indicates that presenting such a duality adequately proves $\theta^{*}$ to be a maximum margin solution.
Recall that there is a “max” operation in Definition 3.2, which makes the swapping of expectation and summation infeasible, meaning that the expected network margin cannot be broken down into the expected margins of individual neurons. To tackle this problem, the class-weighted margin is proposed, whose intuition is similar to label smoothing. Let $\tau:D_{p}\rightarrow\Delta(\mathcal{Y})$ allocate weights to incorrect labels for every data point. Given $(x,y)$ in $D_{p}$ and for any $y^{\prime}\in\mathcal{Y}$, we have $\tau(x,y)[y^{\prime}]\geq 0$ and $\sum_{y^{\prime}\in\mathcal{Y}\backslash\{y\}}\tau(x,y)[y^{\prime}]=1$. Then, we denote a proxy $g^{\prime}$ as the following to solve the issue.
@@
$~{}\underset{(x,y)}{\operatorname*{{\mathbb{E}}}}[g^{\prime}(\theta,x,y)]=~{}\sum_{i=1}^{m}\underset{(x,y)}{\operatorname*{{\mathbb{E}}}}[\phi(\theta_{i},x)[y]-\sum_{y^{\prime}\in\mathcal{Y}\backslash\{y\}}\tau(x,y)[y^{\prime}]\phi(\theta_{i},x)[y^{\prime}]],$
where we can move the summation $\sum_{i=1}^{m}$ out of the expectation $\operatorname*{{\mathbb{E}}}[]$.
3.3 Connection between Training and the Maximum Margin Solutions
Figure 1: Cosine shape of the trained embeddings (hidden layer weights) and corresponding power of Fourier spectrum. The two-layer network with $m=2944$ neurons is trained on $k=4$-sum mod-$p=47$ addition dataset. We even split the whole datasets ($p^{k}=47^{4}$ data points) into the training and test datasets. Every row represents a random neuron from the network. The left figure shows the final trained embeddings, with red dots indicating the true weight values, and the pale blue interpolation is achieved by identifying the function that shares the same Fourier spectrum. The right figure shows their Fourier power spectrum. The results in these figures are consistent with our analysis statements in Lemma 4.2. See Figure 57 in Appendix I.2 for similar results when $k$ is 3 or 5.